Someone help me figure out this problem please my teacher wants to know x1 and x 2 how do i find them and what are day please give me some steps

Solve the equation x^2 - 8 x - 20=0 by completing the square.

Answer by nerdybill(7384)

x^2 - 8x - 20=0
x^2 - 8x = 20
take half 'b', square it, and add to both sides:
x^2 - 8x + 16 = 20 + 16
(x-4)(x-4) = 36
(x-4)^2 = 36
take the square root of both sides:
x-4 = +-6
x = 4+-6
x = {-2, 10}